# Extended Resolution Clause Learning via Dual Implication Points

### v2

@@ Abstract.
Albert Oliveras is supported by grant PID2021-122830OB-C43, funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF: A way of making Europe” and by Barcelogic through research grant C-11423.
@@ 1. Introduction
Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [Kro21], security [XA05], and AI [KS92, Kau06]. As solvers continue to be adopted in increasing complex settings, the demand for greater efficiency and reasoning power by users continues unabated.
While developers continue to improve CDCL SAT solvers, it is simultaneously true that the base solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [AFT11a, PD10], and therefore are fundamentally limited. Hence, solver developers have been actively researching novel algorithms that implement stronger proof systems that go beyond Res. Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [HKB19, OLW+23], bounded variable addition (BVA) [MHB12], symmetry breaking [Sak21], and extended resolution (ER) solvers such as GlucosER [AKS10].
@@
Implementation of the ERCL algorithm requires several additional methods. First, we need a method to identify DIPs, i.e., a technique that takes as input an implication graph and outputs a DIP and does so in time linear in the size of the input. Second, we need a technique replaces this DIP pair with a new variable and appropriately modifies the clause learning algorithm to learn new clauses involving DIPs. Third, we need an ER framework, built on top of a CDCL SAT solver, that enables new variable addition, ER clause addition and deletion, etc. Finally, we need heuristics that specialize the above mentioned methods in a variety of ways, such as clause learning and clause deletion policies that are based on different kinds of DIPs. We implement all of these methods as part of the MapleLCM solver [LLX+17], and refer to the resulting solver as xMapleLCM. In fact our proposed ERCL method, and its implementation, is very general and easily extensible thus encouraging future exploration and specialization efforts with a variety of heuristics.
@@
    DIP: We introduce the concept of dual implication points (DIP), a generalization of UIPs in conflict graphs. We also came up with an algorithm that computes them in linear time. However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [BGO24].
@@
    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random $k$-xor, Tseitin, and Interval Matching, and compare xMapleLCM against leading solvers such as Kissat [BFFH20], CryptoMiniSat [SNC09], SBVA-CaDiCaL [HGH23], and the extended-resolution solver GlucosER [AS09]. Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucosER excel on them. Hence, we can now finally state that extended resolution can be added to CDCL and improve the performance of these solvers, and that there are at least two completely different ways to achieve this. Moreover, a simple heuristic dynamically allows us to detect whether extended resolution is being helpful and turn back to standard CDCL in order to get the best of both worlds. This technique enables xMapleLCM to perform similarly to MapleLCM on the SAT Competition 2023 Main Track instances.
@@ 2. Related Work
The idea of using ER in SAT solving has been studied in various forms in the literature for nearly two decades. The closest approach to ours is GlucosER [AKS10], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$. Hence, information across different conflict analysis steps is used to detect which extended variables to introduce.
@@
More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [MHB12] and its structured version SBVA [HGH23], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses. Essentially, the ultimate goal of BVA techniques is to reduce the size of the formula by reducing the number of clauses at the cost of adding a new variable. One big difference with our work is that this process is done only in a preprocessing step. Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [BH23].
Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [HKB17, HKB20], either via preprocessing steps [RHB22] or via the use of the SDCL algorithm [HKSB17, OLW+23]. While these methods implement proof systems that are stronger than Res, many of them (without new variable addition) are known to be weaker than ER.
@@ 3. Preliminaries
We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [MLM21]. Below we focus on conflict analysis, which is the most relevant ingredient from the CDCL algorithm for this paper. We do so by means of the following example.
@@
$\begin{array}[]{llllll}(1)&y_{1}\lor\lnot x_{1}\lor x_{2}&(6)&\lnot x_{5}\lor x_{7}&(11)&\lnot x_{11}\lor x_{12}\\ (2)&\lnot x_{1}\lor\lnot x_{3}&(7)&x_{6}\lor\lnot x_{7}\lor x_{8}&(12)&x_{10}\lor\lnot x_{11}\lor x_{13}\\ (3)&y_{2}\lor\lnot x_{1}\lor x_{4}&(8)&\lnot y_{3}\lor\lnot y_{4}\lor\lnot x_{5}\lor\lnot x_{9}&(13)&x_{12}\lor\lnot x_{13}\\ (4)&\lnot y_{3}\lor\lnot x_{2}\lor x_{3}\lor\lnot x_{4}\lor x_{5}&(9)&\lnot y_{4}\lor x_{9}\lor\lnot x_{10}\\ (5)&y_{1}\lor\lnot x_{5}\lor\lnot x_{6}&(10)&\lnot y_{5}\lor y_{6}\lor\lnot x_{8}\lor x_{9}\lor x_{11}\\ \end{array}$
@@
The graph clearly shows that if we set $x_{1}$ to true, together with the literals of previous decision levels (the $y$’s), we obtain a conflict. However, the same happens with $x_{5}$, since any path from $x_{1}$ to the conflict necessarily goes through $x_{5}$. Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$. Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [ZMMM01]. It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict. Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$. The quality of a lemma can be assessed by its Literal Block Distance (LBD) [AS09]: the number of different decision levels of the literals in the lemma. The lower the LBD, the better the lemma. In our case, if we are at decision level 5, $y_{4}$ and $y_{6}$ belong to decision level 2, and $y_{5}$ to decision level 4, the LBD of the lemma is 3. $\Box$
@@
Resolution. Given two clauses $l\lor C$ and $\lnot l\lor D$, the resolution inference rule allows one to derive the logical consequence $C\lor D$. It is well known that the lemma derived in Example 3 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment. In fact, it has been proved [PD11, AFT11b] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [Hak85], or Tseitin formulas [Urq87]) are also hard for CDCL SAT solvers.
Extended Resolution (ER). Given two literals $l_{1}$ and $l_{2}$, the extended resolution [Tse83] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$. ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [Coo76]. Incorporating ER to CDCL solvers could potentially enable them to solve such formulas in polynomial time. However, we lack good methods to incorporate ER into CDCL proof search. This is precisely the aim of this paper, namely, incorporating a restricted version of ER into CDCL.
@@ 4. Dual Implication Points
$\overline{x_{12}},x_{13}$ $z\leftrightarrow(\overline{x_{12}}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$x_{11},x_{13}$ $z\leftrightarrow(x_{11}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$\overline{x_{10}},x_{11}$ $z\leftrightarrow(\overline{x_{10}}\land x_{11})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
@@
Figure 2 also shows how a DIP pair can be used to introduce a variable $z$ via extension, and the associated clauses that can be learned. For example, in the third line, the new extension variable $z$ is introduced with the three clauses $\lnot z\lor\lnot x_{10}$, $\lnot z\lor x_{11}$ and $x_{10}\lor\lnot x_{11}\lor\lnot z$ which express the condition $z\leftrightarrow(\lnot x_{10}\land x_{11})$. From the conflict graph, this allows inferring the clauses $\lnot z$ (the “post-DIP” learned clause) and $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$ (the “pre-DIP” learned clause). Since the post-DIP learned clause does not have any variables from lower decision levels, we can also infer the 2-clause $x_{10}\lor\lnot x_{11}$ either instead of or in addition to introducing $z$ and the pre- and post-DIP clauses. Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [Bac02] or the recent work of Biere et al. [BFW23] or Buss et al. [BKW24].
In general, for any literals $a,b$ that form a DIP in the above fashion, we may introduce an extension variable $z\leftrightarrow a\land b$, and learn (1) a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ (i.e. $f\land C\rightarrow z$), where $f$ is the first UIP and $C$ the set of literals from previous decision levels that have an edge in the conflict graph to any literal appearing after the first UIP and no later than the DIP pair; and (2) a post-DIP clause of the form $\lnot z\lor\lnot D$ (i.e. $z\land D\rightarrow\perp$), where $D$ contains those literals from previous decision levels with an edge to any literal appearing strictly after the DIP pair. The correctness of adding these two clauses is guaranteed because they can be generated via a sequence of resolution steps on the formula and the clauses that define the extended variable $z$.
For example, the last line of the table in Figure 2 shows the case $z\leftrightarrow(x_{8}\land\overline{x_{9}})$, where $f$ is $x_{5}$, and $C$ and $D$ are $\lnot y_{1}\land y_{3}\land y_{4}$ and $y_{4}\land y_{5}\land\lnot y_{6}$, respectively. In Section 5.2 we discuss many possible ways that DIP extension variables and clauses may be introduced into CDCL solvers. The rest of this section describes how to find DIPs.
@@ 4.1. An Algorithm for Finding DIPs
A conventional CDCL algorithm maintains a trail of the literals set true, in the order they were set, and this allows finding the UIP very quickly. Finding the DIPs is much more complex than finding the UIP, and requires several traversals of the conflict graph between the first-UIP and the conflict node. Nonetheless, it is possible to find all DIPs very quickly, even in linear time.
@@
We recast the DIP-finding problem in terms of a general directed graph $G$. Let $G=(V,E)$ be a directed graph with two distinguished vertices $s$ and $t$. We assume that $s$ is 2-connected to $t$ in that there is a pair of vertex-disjoint paths $\pi_{1}$ and $\pi_{2}$ from $s$ to $t$ that share no vertices apart from $s$ and $t$. By the vertex-version of Menger’s theorem [Men27], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$. Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Dominator (TVD).
@@
Our goal is to find all possible TVDs efficiently and in linear time. An algorithm for this is discussed in our companion paper [BGO24]; for space reasons, we give only an abbreviated discussion of the algorithm here. The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path construction to find two vertex-disjoint paths from $s$ to $t$. An example is shown in Figure 3, where the two paths are $a_{0},\ldots,a_{\ell}$ and $b_{0},\ldots,b_{k}$ where $a_{0}=b_{0}=s$ and $a_{\ell}=b_{k}=t$. These paths are called $\pi_{a}$ and $\pi_{b}$, respectively. Henceforth, a path is a directed path without any repeated nodes. The internal vertices of a path $\pi$ are the vertices on $\pi$ other than the first and last vertices. Two paths are said to be vertex-disjoint if they have no internal vertices in common. A path $\pi$ avoids $\pi_{a}$ and $\pi_{b}$ if it is vertex-disjoint from both paths.
$s$$a_{1}{=}a^{\prime}_{1}$$a_{2}{=}a^{\prime}_{2}$$a_{3}{=}a^{\prime}_{3}$$a_{4}$$a_{5}{=}a^{\prime}_{4}$$b_{1}{=}b^{\prime}_{1}$$b_{2}{=}b^{\prime}_{2}$$b_{3}$$b_{4}$$b_{5}{=}b^{\prime}_{3}$$t$$c$$d$
@@
[BGO24] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$. A node $b_{i}$ being bypassed is defined similarly.
@@
[BGO24] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex dominator (TVD) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
Theorem 1 is proved in [BGO24]. The theorem holds for both acyclic and cyclic directed graphs; however, for our applications to CDCL we are interested only in acyclic graphs since the conflict graph is always acyclic.
@@
Theorem 1 is used in [BGO24] to give an efficient, linear time algorithm for finding all TVDs. The algorithm has five phases. The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVDs. Full details are given [BGO24], which is available online in preprint form. In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVDs is negligible.

### v3

@@ Abstract.
Albert Oliveras is supported by grant PID2021-122830OB-C43, funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF: A way of making Europe”; by grant PID2024-157044OB-C32, funded by MICIU /AEI /10.13039/501100011033 / FEDER, UE. and by Barcelogic through research grant C-11423.
@@ 1. Introduction
Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [HandbookSATSoftwareVerification], security [xie2005security], and AI [satplan1, satplan2]. As solvers continue to be adopted in increasing complex settings, the demand for greater efficiency and reasoning power by users continues unabated.
While developers continue to improve CDCL SAT solvers, it is simultaneously true that the base solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [AFT11ClauseLearning, PipatsrisawatDarwiche:modernclauselearning], and therefore are fundamentally limited. Hence, solver developers have been actively researching novel algorithms that implement stronger proof systems that go beyond Res. Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [HKB:RedundancySDCL, SDCLMaxSAT], bounded variable addition (BVA) [BVA], symmetry breaking [HandbookSATSymmetry], and extended resolution (ER) solvers such as GlucosER [Audemard2010ARO].
@@
Implementation of the ERCL algorithm requires several additional methods. First, we need a method to identify DIPs, i.e., a technique that takes as input an implication graph and outputs a DIP and does so in time linear in the size of the input. Second, we need a technique replaces this DIP pair with a new variable and appropriately modifies the clause learning algorithm to learn new clauses involving DIPs. Third, we need an ER framework, built on top of a CDCL SAT solver, that enables new variable addition, ER clause addition and deletion, etc. Finally, we need heuristics that specialize the above mentioned methods in a variety of ways, such as clause learning and clause deletion policies that are based on different kinds of DIPs. We implement all of these methods as part of the MapleLCM solver [MapleLCM], and refer to the resulting solver as xMapleLCM. In fact our proposed ERCL method, and its implementation, is very general and easily extensible thus encouraging future exploration and specialization efforts with a variety of heuristics.
@@
    DIP: We introduce the concept of dual implication points (DIP), a generalization of UIPs in conflict graphs. We also came up with an algorithm that computes them in linear time. However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [BGO:TVBs].
@@
    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random $k$-xor, Tseitin, and Interval Matching, and compare xMapleLCM against leading solvers such as Kissat [kissat], CryptoMiniSat [cryptominisat], SBVA-CaDiCaL [SBVA], and the extended-resolution solver GlucosER [LBDAudemardSimon]. Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucosER present much better performance. This demostrates that extended resolution can be added to CDCL and improve the performance of these solvers; GlucosER and our DIP-based methods give two different ways to achieve this. Moreover, for our DIP-based learning, a simple heuristic allows us to detect on-the-fly whether extended resolution is being helpful and revert to using standard CDCL in order to get the best of both worlds. This technique enables xMapleLCM to perform similarly to MapleLCM on the SAT Competition 2023 Main Track instances.
@@ 2. Related Work
The idea of using ER in SAT solving has been studied in various forms in the literature for nearly two decades. The closest approach to ours is GlucosER [Audemard2010ARO], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$. Hence, information across different conflict analysis steps is used to detect which extended variables to introduce.
@@
More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [BVA] and its structured version SBVA [SBVA], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses. Essentially, the ultimate goal of BVA techniques is to reduce the size of the formula by reducing the number of clauses at the cost of adding a new variable. One big difference with our work is that this process is done only in a preprocessing step. Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [ExtendedResolutionBDDs].
Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [ShortProofs, StrongExtensionFree], either via preprocessing steps [preprocessingPR] or via the use of the SDCL algorithm [PruningThroughSatisfaction, SDCLMaxSAT]. While these methods implement proof systems that are stronger than Res, many of them (without new variable addition) are known to be weaker than ER.
Another related work is [PD08Learning], where the authors introduce a new conflict analysis procedure that learns lemmas that contain two literals of the last decision level. These two literals are indeed a DIP. However, there can be many other candidate DIP’s that are not detected by the methods of [PD08Learning]. Furthermore, as we will show in Section 4.1, their procedure might sometimes fail to detect any DIPs at all. An additional, important differences w.r.t. our work is that [PD08Learning] do not introduce a new variable representing the conjunction of the DIP members. Hence, this is not an extended resolution-based approach and, from the proof-complexity point of view, is not more powerful than standard CDCL or general resolution.
@@ 3. Preliminaries
We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [HandbookCDCL]. Below we focus on conflict analysis, which is the most relevant ingredient from the CDCL algorithm for this paper. We do so by means of the following example.
@@
[…]
@@
The graph clearly shows that if we set $x_{1}$ to true, together with the literals of previous decision levels (the $y$’s), we obtain a conflict. However, the same happens with $x_{5}$, since any path from $x_{1}$ to the conflict necessarily goes through $x_{5}$. Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$. Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [ConflictAnalysis]. It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict. Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$. The quality of a lemma can be assessed by its Literal Block Distance (LBD) [LBDAudemardSimon]: the number of different decision levels of the literals in the lemma. The lower the LBD, the better the lemma. In our case, if we are at decision level 5, $y_{4}$ and $y_{6}$ belong to decision level 2, and $y_{5}$ to decision level 4, the LBD of the lemma is 3. $\Box$
Resolution. Given two clauses $l\lor C$ and $\lnot l\lor D$, the resolution inference rule allows one to derive the logical consequence $C\lor D$. It is well known that the lemma derived in Example 3 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment. In fact, it has been proved [CDCLResolutionKnot, CDCLResolutionAtserias] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [haken1985intractability], or Tseitin formulas [TseitinHardForResolution]) are also hard for CDCL SAT solvers.
Extended Resolution (ER). Given two literals $l_{1}$ and $l_{2}$, the extended resolution [tseitin1983complexity] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$. ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [PigeonholeExtendedResolution]. Incorporating ER to CDCL solvers could potentially enable them to solve such formulas in polynomial time. However, we lack good methods to incorporate ER into CDCL proof search. This is precisely the aim of this paper, namely, incorporating a restricted version of ER into CDCL.
@@ 4. Dual Implication Points
$\overline{x_{12}},x_{13}$ $\leftrightarrow(\overline{x_{12}}\land x_{13}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$x_{11},x_{13}$ $\leftrightarrow(x_{11}\land x_{13}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$\overline{x_{10}},x_{11}$ $\leftrightarrow(\overline{x_{10}}\land x_{11}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
@@
Figure 2 also shows how a DIP pair can be used to introduce a variable $z$ via extension, and the associated clauses that can be learned. For example, in the third line, the new extension variable $z$ is introduced with the three clauses $\lnot z\lor\lnot x_{10}$, $\lnot z\lor x_{11}$ and $x_{10}\lor\lnot x_{11}\lor\lnot z$ which express the condition $z\leftrightarrow(\lnot x_{10}\land x_{11})$. From the conflict graph, this allows inferring the clauses $\lnot z$ (the “post-DIP” learned clause) and $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$ (the “pre-DIP” learned clause), both of which will be formally defined in the next paragraph. Since the post-DIP learned clause does not have any variables from lower decision levels, we can also infer the 2-clause $x_{10}\lor\lnot x_{11}$ either instead of or in addition to introducing $z$ and the pre- and post-DIP clauses. Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [Bacchus:EnhancingDP] or the recent work of Biere et al. [BFW:CadiBack] or Buss et al. [BKVW:DualDFS].
In general, for any literals $a,b$ that form a DIP in the above fashion, we may introduce an extension variable $z\leftrightarrow a\land b$, and learn (1) a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ (i.e. $f\land C\rightarrow z$), where $f$ is the first UIP and $C$ the set of literals from previous decision levels that have an edge in the conflict graph to any literal appearing after the first UIP and no later than the DIP pair; and (2) a post-DIP clause of the form $\lnot z\lor\lnot D$ (i.e. $z\land D\rightarrow\perp$), where $D$ contains those literals from previous decision levels with an edge to any literal appearing strictly after the DIP pair. The correctness of adding these two clauses is guaranteed because they can be generated via a sequence of resolution steps on the formula and the clauses that define the extended variable $z$. We can also prove their correctness with a more conflict-graph oriented argument: it should be obvious from the conflict graph that if we assert the 1UIP $f$ and all literals in $C$, we can obtain $a$ and $b$ via unit propagation. This means that $f\land C\rightarrow a\land b$ is a valid formula which, after replacing $a\land b$ by their definition $z$, becomes the pre-DIP clause $\lnot f\lor\lnot C\lor z$. Similarly, if we assert $a,b$ and all literals in $D$, we get a conflict by unit propagation. This means that $\lnot a\lor\lnot b\lor\lnot D$ is a valid clause. If we apply two resolution steps with the clauses $a\lor\lnot z$ and $b\lor\lnot z$ we obtain the post-DIP clause $\lnot z\lor\lnot D$.
For example, the last line of the table in Figure 2 shows the case $z\leftrightarrow(x_{8}\land\overline{x_{9}})$, where $f$ is $x_{5}$, and $C$ and $D$ are $\lnot y_{1}\land y_{3}\land y_{4}$ and $y_{4}\land y_{5}\land\lnot y_{6}$, respectively. In Section 5.2 we discuss many possible ways that DIP extension variables and clauses may be introduced into CDCL solvers. The rest of this section describes how to find DIPs and its content is not strictly necessary to understand the rest of the paper.
@@ 4.1. An Algorithm for Finding DIPs
$x_{1}$$x_{3}$$x_{4}$$x_{2}$$x_{5}$$x_{6}$$x_{7}\perp$
 Trail Derived clause $\perp$ $\lnot x_{5}\lor\lnot x_{6}\lor\lnot x_{7}$ $x_{5}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{6}\lor\lnot x_{7}$ $x_{7}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{6}\lor\lnot x_{2}$ $x_{6}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{2}$ $x_{4}$ $\lnot x_{3}\lor\lnot x_{1}\lor\lnot x_{2}$ $x_{3}$ $\lnot x_{1}\lor\lnot x_{2}$ $x_{2}$ $\lnot x_{1}$ $x_{1}$
Figure 3. Conflict graph associated with Example 3, with the literals in blue being the only DIP. Next to it, the resulting trail and the sequence of intermediate clauses obtained if standard 1UIP learning is applied.
A conventional CDCL algorithm maintains a trail of the literals set true, in the order they were set, and this allows finding the UIP very quickly. The idea is initialize a clause $C$ as the conflicting clause and, following the order of the trail from the most recent literal $l$ backwards, apply resolution between $C$ and the reason of $l$ to get a new clause $C$. This is repeated until $C$ only has only one literal of the current decision level.
One could think that finding DIPs can be done by applying the same algorithm and stopping when $C$ has exactly two literals of the current decision level. This is what is done in [PD08Learning] but, as we now show, this procedure is not complete for detecting DIPs. {exa} Consider the set of clauses $\{\lnot x_{1}\lor x_{2},\allowbreak\;\lnot x_{1}\lor x_{3},\allowbreak\;\lnot x_{1}\lor x_{4},\allowbreak\;\lnot x_{3}\lor\lnot x_{4}\lor x_{5},\allowbreak\;\lnot x_{2}\lor x_{6},\allowbreak\;\lnot x_{2}\lor x_{7},\allowbreak\;\lnot x_{5}\lor\lnot x_{6}\lor\lnot x_{7}\}$. Setting decision $x_{1}$ and applying unit propagation finds a conflict. This is illustrated in the conflict graph of Figure 3, where the only DIP is the pair $\{x_{2},x_{5}\}$. In the same figure, we can see the corresponding trail and the intermediate clauses obtained in a standard conflict analysis procedure. Note that the procedure is unable to detect the DIP.
@@
Finding the DIPs is much more complex than finding the UIP, and requires several traversals of the conflict graph between the first-UIP and the conflict node. Nonetheless, it is possible to find all DIPs very quickly, even in linear time.
@@
We recast the DIP-finding problem in terms of a general directed graph $G$. Let $G=(V,E)$ be a directed graph with two distinguished vertices $s$ and $t$. We assume that $s$ is 2-connected to $t$ in that there is a pair of vertex-disjoint paths $\pi_{1}$ and $\pi_{2}$ from $s$ to $t$ that share no vertices apart from $s$ and $t$. By the vertex-version of Menger’s theorem [Menger:Kurventheorie], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$. Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Dominator (TVD).
@@
Our goal is to find all possible TVDs efficiently and in linear time. An algorithm for this is discussed in detail in our companion paper [BGO:TVBs]; for space reasons, we give only an abbreviated discussion of the algorithm here and direct the reader who wants a deep understanding of the algorithm to [BGO:TVBs]. The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path [^3] This can also be computed in linear time by applying Ford-Fulkerson’s algorithm to a small modification of the conflict graph in order look for a flow of 2. construction to find two vertex-disjoint paths from $s$ to $t$. An example is shown in Figure 4, where the two paths are $a_{0},\ldots,a_{\ell}$ and $b_{0},\ldots,b_{k}$ where $a_{0}=b_{0}=s$ and $a_{\ell}=b_{k}=t$. These paths are called $\pi_{a}$ and $\pi_{b}$, respectively. Henceforth, a path is a directed path without any repeated nodes. The internal vertices of a path $\pi$ are the vertices on $\pi$ other than the first and last vertices. Two paths are said to be vertex-disjoint if they have no internal vertices in common. A path $\pi$ avoids $\pi_{a}$ and $\pi_{b}$ if it is vertex-disjoint from both paths.
[…]
@@
[BGO:TVBs] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$. A node $b_{i}$ being bypassed is defined similarly.
@@
[BGO:TVBs] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex dominator (TVD) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
Theorem 1 is proved in [BGO:TVBs]. The theorem holds for both acyclic and cyclic directed graphs; however, for our applications to CDCL we are interested only in acyclic graphs since the conflict graph is always acyclic.
@@
Theorem 1 is used in [BGO:TVBs] to give an efficient, linear time algorithm for finding all TVDs. The algorithm has five phases. The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVDs. Full details are given [BGO:TVBs], which is available online in preprint form. In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVDs is negligible.